OpenStudy (anonymous):

Increasing a number by 5 and multiplying the result by 7 is the same as multiplying the number by -5 and subtracting 1 from the result. What is the number?

OpenStudy (superdavesuper):

lets assume the number is x okay? so plz write out "Increasing a number by 5 and multiplying the result by 7" in term of x

OpenStudy (anonymous):

(x+5)*7?

OpenStudy (superdavesuper):

VERY GOOD :) would u do the same for "multiplying the number by -5 and subtracting 1 from the result" plz?

OpenStudy (anonymous):

(-5x)-1?

OpenStudy (superdavesuper):

RIGHT again hahahaaa u hardly need my help :P So ur Q said the two are the same right?

OpenStudy (anonymous):

yes. do i equal them to eachother?

OpenStudy (superdavesuper):

yup yup show me wat u got plz

OpenStudy (anonymous):

(x+5)*7=(-5x)-1 7x+35=5x?

OpenStudy (superdavesuper):

OK on the left....no simplification on the right. Then u just move all x terms to one side and everything else to the other side of =

OpenStudy (superdavesuper):

7x+35=-5x-1 go on from here plz :)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

12x=-36?

OpenStudy (anonymous):

x=-3?

OpenStudy (superdavesuper):

right again :)

OpenStudy (anonymous):

-3 is the number?

OpenStudy (superdavesuper):

something wrong w that answer? yes, -3 is the number!

OpenStudy (anonymous):

Thank you :)

OpenStudy (superdavesuper):

welcome n thanx for the medal :)

OpenStudy (anonymous):

Sure thing :)